OpenStudy (anonymous):

todd is planning an graduation party at the house of bratwurst. It costs $150.00 to rent a tent plus $11.50 per person for food. If todd can spend at most $450.00, find the greatest number of people he can invite to the party?

OpenStudy (mathstudent55):

Can you write an expression that shows how much the party costs. Let y = cost, and x = number of people. Can you write an expression for y in terms of x?

OpenStudy (anonymous):

so for y would that be 450 or would it be 150?

OpenStudy (mathstudent55):

No. I'll explain. cost = cost of tent + cost for people

OpenStudy (mathstudent55):

cost of tent = 150 So, so far we have y = 150 + cost of people Ok?

OpenStudy (anonymous):

150 + 11.50x = 450 solve for x

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

The cost for the people is according to this pattern: For 1 person it costs $11.50 For 2 people is costs 2 * $11.50 For 3 people it costs 3 * $11.50 So, For x people it costs x * $11.50 Ok?

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

x * $11.50 can be written as 11.50x

OpenStudy (mathstudent55):

Now we can write the expression for th ecost of the party including the tent and the people: cost of party = cost of tent + cost of people y = 150 + 11.50x

OpenStudy (mathstudent55):

ok?

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

We are told Todd can spend at most $450. The maximum cost he can afford is $450. We let y equal the cost of the party, so we can write: 450 = 150 + 11.50x

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

Now you solve the equation for x. That gives you the grwatest number of people he can invite to the party. If you get a number with a decimal part, like 23.8, then you have to round down to 23.

OpenStudy (anonymous):

got it

OpenStudy (mathstudent55):

great